# Emotional Dimension Control in Language Model-Based Text-to-Speech: Spanning a Broad Spectrum of Human Emotions

### v1

@@ III-C Objective Evaluation
Figure 5: XAB test to assess the emotion intelligibility for 4 emotion pairs (‘Surprise’ and ‘Alert’, ‘Relaxed’ and ‘Protected’, ‘Elation’ and ‘Pleasure’, ‘Anger’ and ‘Anxiety’).
III-D Subjective Evaluation
@@
We conducted listening experiments, with 10 native English speakers each listening to a total of 260 synthesized speech samples. We first evaluate our proposed framework on the zero-shot TTS task, where the ED is predicted based on the prompt speech (‘Proposed w/ ED Prediction’). We selected ‘Phonetic-Enhanced VALL-E’ [24] as the baseline. Compared to the popular VALL-E model, this baseline offers greater robustness and enhanced zero-shot capabilities [24]. A mean opinion score (MOS) test is conducted to evaluate the speech naturalness, where participants were asked to score each sample on a 5-point scale (5: Excellent, 4: Good, 3: Satisfactory, 2: Unsatisfactory, 1: Bad). As shown in Fig. 4(a), our proposed with ED prediction achieves a score of 3.67, outperforming the baseline score of 2.87.
We additionally conducted two AB preference tests, where participants were asked to choose the better speech sample in terms of speech naturalness and prosody similarity with the prompt speech, respectively. As shown in Fig. 4(b), the proposed framework with ED prediction consistently outperforms the baseline by achieving a higher preference percentage 78.9%. These results demonstrate that our proposed method significantly enhances speech naturalness and prosody consistency with the prompt by using ED to guide acoustic prediction during TTS training.
Next, we perform emotion perception experiments. Note that participants were not provided with explicit emotional references and they were asked to make judgments based on their innate emotional concepts. This approach aligns with methodologies used in emotion perception studies [45] and speech synthesis evaluation [31]. We first conducted a Comparative MOS (CMOS) test towards basic emotions1. In the proposed framework, the ED was manually assigned to simulate these emotions. In this test, participants were presented with pairs of synthesized neutral and emotional speech and asked to rate each sample on a 5-point scale based on how easily they could identify the intended emotion (5: Very Easy, 4: Easy, 3: Fair, 2: Difficult, 1: Very Difficult). The final results in Table II are measured as the comparative average score (CMOS) of emotional speech relative to the neutral counterpart. We observe that all emotions consistently achieve positive scores, indicating that listeners could perceive the intended emotions over the neutral one. Notably, ‘Surprise’ achieved the highest CMOS 0.92, demonstrating greater intelligibility, while ‘Sad’ had the lowest score 0.15, indicating it is closer to a neutral emotion.
@@
We further conducted four XAB tests, grouping 4 emotion pairs 2 : 1) ‘Alert’ (0.5, 0.5, 0.5) vs. ‘Surprise’ (0.5, 0.5, -0.1); 3) ‘Relaxed’ (0.5, -0.5, 0.2) vs. ‘Protected’ (0.5, -0.2, -0.5); 3) ‘Anger’ (-1, 1, 1) vs. ‘Anxiety’ (-1, 1, -1); 4) ‘Elation’ (1, 1, 0.5) vs. ‘Pleasure’ (1, 0, 0.5). For each pair, participants were asked to correctly match the emotions, such as choosing between option A: X is Anger and Y is Anxiety and option B: X is Anxiety and Y is Anger. The results are summarized in Fig. 5. For all 4 emotion pairs, the correct match consistently exceeds 50%, with the ‘Elation vs. Pleasure’ pair achieving the highest accuracy at around 71.1%. The results in Fig.5 demonstrate a positive alignment with 2 key findings in [23]:
@@
  • ‘Dominance’ is a key factor in distinguishing ‘Angry’ from ‘Anxious’, ‘Alert’ from ‘Surprised’, and ‘Relaxed’ from ‘Protected’ [^7] The first word in each pair involves dominance; the second word involves submissiveness.;
  • ‘Pleasure’ and ‘Arousal’ combine to contribute to ‘Elation’;
@@
IV Conclusion
Inspired by the success of language models in a variety of generative tasks, much of current research prioritizes data collection, aiming to expand training data to push the boundaries of the state-of-the-art. In contrast, this paper explores how earlier studies on emotion theory can can advance research in emotional TTS. We introduce a TTS framework capable of controlling 3 emotional dimensions—pleasure, arousal, and dominance—during runtime. This framework not only synthesizes a wide range of emotional styles without the need for any emotional speech data for TTS training, but also enhances naturalness and prosody consistency with the speech prompt in zero-shot scenarios.

### v2

@@ 3.3 Objective Evaluation
Figure 4: XAB test to assess the emotion intelligibility for 4 emotion pairs (‘Surprise’ and ‘Alert’, ‘Relaxed’ and ‘Protected’, ‘Excited’ and ‘Pleasure’, ‘Angry’ and ‘Anxious’). ED values were manually assigned based on Table 1 to simulate different emotions (‘ED Control’).
3.4 Subjective Evaluation
@@
We conducted listening experiments, with 10 native English speakers each listening to a total of 260 synthesized speech samples. We first evaluate our proposed framework on the zero-shot TTS task, where the ED is predicted based on the prompt speech (‘Proposed’). We selected ‘CosyVoice’ [13] as the baseline. A mean opinion score (MOS) test is conducted to evaluate the speech naturalness, where participants were asked to score each sample on a 5-point scale (5: Excellent, 4: Good, 3: Satisfactory, 2: Unsatisfactory, 1: Bad). As shown in Table 2, our proposed framework with ED prediction achieves a score of 4.54, outperforming the baseline score of 4.36, while lower than the ground-truth score of 4.80. It shows that our proposed framework with ED prediction achieves better naturalness than the baseline CosyVoice.
We additionally conducted a mean opinion score to assess emotional intelligibility (E-MOS), where participants were asked to rate each speech sample in terms of emotion perception. We evaluated our proposed framework against two baselines: MixedEmotion [25] and CosyVoice [13]. As shown in Fig.3, our proposed framework with ED prediction consistently outperforms CosyVoice with a higher E-MOS score, while MixedEmotion, which trains on in-domain emotional data, achieves the top score below the ground truth. In contrast, both CosyVoice and our proposed system rely solely on zero-shot emotion cloning. These results indicate that our proposed method effectively captures and replicates emotional nuances in synthesized speech, thereby enhancing emotion cloning performance.
@@
We further conducted four XAB tests, grouping 4 emotion pairs 2 : 1) ‘Anger’ vs. ‘Anxious’; 2) ‘Excited’ vs. ‘Pleasure’; 3) ‘Protected’ vs. ‘Relaxed’ ; 4) ‘Alert’ vs. ‘Surprise’. For each pair, participants were asked to correctly match the emotions, such as choosing between option A: X is Anger and Y is Anxiety and option B: X is Anxiety and Y is Anger. The results are summarized in Fig. 4. For all 4 emotion pairs, the correct match consistently exceeds 50%, with the ‘Anger vs. Anxious’ pair achieving the highest accuracy at around 83.75%. The results in Fig.4 demonstrate a positive alignment with 2 key findings in [18]:
@@
  • ‘Dominance’ is a key factor in distinguishing ‘Angry’ from ‘Anxious’, ‘Alert’ from ‘Surprised’, and ‘Relaxed’ from ‘Protected’ [^4] The first emotion in each pair involves dominance; the second emotion involves submissiveness.;
  • ‘Pleasure’ and ‘Arousal’ combine to contribute to ‘Excitement’;
@@
4 Conclusion
In this paper, we proposed an emotional TTS framework that integrates established emotion theory with language model-based speech generation. By controlling pleasure, arousal, and dominance, our proposed system synthesizes a wide range of emotional speech without relying on large-scale emotion-labeled data, while preserving naturalness in zero-shot scenarios. This work highlights the value of combining psychological insights with data-driven modeling. Future directions include exploring dynamic within-utterance emotion control and multilingual adaptation, paving the way for more versatile, affect-aware TTS solutions.